OpenStudy (anonymous):

What is the maximum vertical distance between the line y = x + 30 and the parabola y = x2 for −5 ≤ x ≤ 6?

OpenStudy (anonymous):

vertical distance implies maximize \(|y_1-y_2|\) where \(y_1=x+30\text{ and }y_2=x^2\)

OpenStudy (anonymous):

I plug in y1 and y2 to |y1-y2|?

OpenStudy (anonymous):

what class are you in? are you taking calculus?

OpenStudy (anonymous):

Yes I'm taking calculus

OpenStudy (anonymous):

so \(x+30 - x^2\) is the function you want to maximize in the interval \(5 \le x \le 6\), right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and \(x+30 \ge x^2\) in the interval \(-5\le x\le 6\) right? so take the derivative set it equal to 0 and see if that x is in your interval. if it is, then that's the max (you should check to make sure it is a max but it will be).

OpenStudy (anonymous):

once you know the value of x, plug it into the distance function \(\left(-x^2+x+30\right)\) and then you'll have the max distance.

OpenStudy (anonymous):

I found my inverse which is: -2x+1 right? And I equal it to zero and I got x=1/2 Then I plugged it in the formula and I got 30.75?

OpenStudy (anonymous):

there you go! except it should be 30.5 - .25 = 30.25